OpenStudy (anonymous):

8x − 2 = y 5y − 12x = -3

OpenStudy (anonymous):

plug y into 2nd equation

OpenStudy (ajprincess):

8x − 2 = y 5y − 12x = -3 5(8x-2)-12x=-3 40x-10-12x=-3 Solve for x. Substitute the value of x in 8x-2=y to find y.